OpenStudy (anonymous):

Coral created a painting that was 4 1/3 ft long and 1 1/5 t wide. What was the area of Coral's painting? Express your answer in simplest form.

zepdrix (zepdrix):

Area is given by length times width.\[\large\rm A=\ell \times w\]So we'll need to multiply these mixed numbers together to find our area,\[\large\rm A=4\frac{1}{3}\times1\frac{1}{5}\]

zepdrix (zepdrix):

This will easier to multiply if we first express our mixed numbers as `improper fractions`.

zepdrix (zepdrix):

Do you remember how to do that?

zepdrix (zepdrix):

Here is how we would deal with the first number:\[\large\rm 4\frac{1}{3}\quad=\frac{4\cdot3+1}{3}\quad= \frac{13}{3}\]

OpenStudy (anonymous):

so wat is the answer

OpenStudy (anonymous):

@tootzrll

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

\[4\frac{1}{3}\times1\frac{1}{5}\]

OpenStudy (tootzrll):

what he said xD

OpenStudy (anonymous):



OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

@tootzrll

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

no